Professor Wendy Smith has been offered the following​ opportunity: A law firm would like to retain her for an upfront payment of $50,000. In​ return, for the next year the firm would have access to eight hours of her time every month. As an alternative payment​ arrangement, the firm would pay Professor​ Smith's hourly rate for the eight hours each month. ​ Smith's rate is $540 per hour and her opportunity cost of capital is 15% per year. What does the IRR rule advise regarding the payment​ arrangement? (Hint: Find the monthly rate that will yield an effective annual rate of 15%​.) What about the NPV​ rule?
Answer #1

Monthly Rate = 8(540) = $4,320

Calculating Present Value of Monthly Payment,

Using TVM Calculation,

PV = [FV = 0, PMT = 4,320, I = 0.15/12, N = 12]

PV = $47,862.63

NPV = -50,000 + 47,862.63

NPV = -$2,137.37